OpenStudy (anonymous):

No one will help me with this question... Can someone please help? I'll give a metal.

OpenStudy (anonymous):

A toy rocket was launched from the ground. The function f(x) = -16x2 + 256x shows the height of the rocket f(x), in feet, from the ground at time x seconds. What is the axis of symmetry of the graph of f(x), and what does it represent? (3 points) x = 8; it takes 8 seconds to reach the maximum height and 8 seconds to fall back to the ground x = 16; it takes 16 seconds to reach the maximum height and 16 seconds to fall back to the ground x = 8; it takes 8 seconds to reach the maximum height and 16 seconds to fall back to the ground x = 16, it takes 16 seconds to reach the maximum height and 32 seconds to fall back to the ground

OpenStudy (dtan5457):

Still need help on this?

OpenStudy (dtan5457):

Ok, the axis of symmetry is measured by -b/2a You know what your b and a values are?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

I'll warn you now. I'm horrible with math.

OpenStudy (dtan5457):

A quadratic's standard form is ax^2+bx+c -16x^2 + 256x There is no c value here, but a and b values are clearly here. Can you tell me what they are?

OpenStudy (anonymous):

wait. wouldnt a- 16 and then b- 256

OpenStudy (dtan5457):

Exactly, so plug into -b/2a -256/-32=?

OpenStudy (anonymous):

256/ -64 =-4

OpenStudy (anonymous):

oh wait. its just 4 because 256 isnt positive

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

B=256 A=-16 -b=-256 2a=-32 -256/-32=?

OpenStudy (anonymous):

8?

OpenStudy (dtan5457):

Yep, so immediately you can cross out B and D from your answer choices

OpenStudy (dtan5457):

Now, since 8 is the axis of symmetry, it takes 8 seconds for the rocket to reach full height. To know what the time it takes for it to fall back to the ground...y must equal 0. The choices are either 8 or 16. Plug 8 and 16 to -16^x2 + 256x, and see which one equals 0.

OpenStudy (anonymous):

I'd plug the two into x right

OpenStudy (dtan5457):

Yep, one by one.

OpenStudy (anonymous):

so itd be c?

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

Yep! 16 would make the y=0.

OpenStudy (anonymous):

thank you!

OpenStudy (dtan5457):

yw